# Stabilizability and lower spectral radius for linear switched systems with singular matrices

### v2

@@ 2.2 Joint spectral subradius and stabilizability radius
In this paper, a matrix set $\mathcal{M}$ is called irreducible if for any $\mathcal{M}_{1}\subset\mathcal{M}$, we have
@@
Not all linear discrete time switching systems are irreducible. However, it is sufficient to only consider the stabilizability radius of irreducible matrix sets. This is because an irreducible subset with the same stabilizability radius exists in any reducible matrix set (Corollary 3.4 in [18]).
@@
The $\tilde{\rho}=\check{\rho}$ if the matrix set is irreducible and contains a singular matrix with one dimensional image.
@@
For this irreducible matrix set, we know that the optimal sequence of $\tilde{\rho}$ contains at least one singular matrix. We can always obtain the optimal sequence of matrices in two steps for such situations. The first step is finding the optimal sequence before applying the singular matrix, and then the second step is finding the optimal sequence starts from the image of the singular matrix. The first part of sequence has no influence on the stabilizability radius, because it doesn’t repeat afterwards. Then it is sufficient to find the optimal sequence only based on vectors in the image of the singular matrix, which have same direction. Consequently, this optimal sequence doesn’t depend on the initial vectors. ∎
Corollary 2.7.
We have $\tilde{\rho}=\check{\rho}$ for any two dimensional irreducible matrix set with at least one singular matrix.
However, for arbitrary matrix set with singular matrices, we don’t have such properties. Consider Example 2.4, we have $\check{\rho}=1$ but this system is pointwise stabilizable as it behaves the same as Example 2.3 (proved in [17]).
3 Two dimensions
@@ 3.1 A singular and a rotation matrix
Inspired by an example in [17], we focus on a two dimensional switched system in this section.
System 1.
@@
Let $\lambda_{1}$ and $\lambda_{2}$, $v_{1}$ and $v_{2}$ denote two eigenvalues and eigenvectors of $M_{1}$, $\lambda_{3}$ and $\overline{\lambda}_{3}$ denote two eigenvalues of $M_{2}$. If $\lambda_{1}=\lambda_{2}=0$, $M_{1}$ is nilpotent, then System 1 becomes trivial. In the following we use $\lambda_{2}\in\mathbb{R}\setminus\{0\}$ and $\lambda_{1}=0$.
The matrix $M_{2}$ is similar to its real Jordan form (see for example Theorem 3.4.1.5. in [8]),
$J=|\lambda_{3}|\left(\begin{matrix}\cos\alpha&\sin\alpha\\ -\sin\alpha&\cos\alpha\end{matrix}\right)$
where $\cos\alpha=(\lambda_{3}+\overline{\lambda}_{3})/2|\lambda_{3}|$. Thus, there exists an invertible matrix $P$ such that $M_{2}=PJP^{-1}$. Let $M^{\prime}_{1}=P^{-1}M_{1}P$, then the system 1 is stabilizable if and only if the following switched system is stabilizable.
System 2.
@@
$M^{\prime}_{1}=\left(\begin{matrix}a^{\prime}_{11}&a^{\prime}_{12}\\ a^{\prime}_{21}&a^{\prime}_{22}\end{matrix}\right),\quad J=|\lambda_{3}|\left(\begin{matrix}\cos\alpha&\sin\alpha\\ -\sin\alpha&\cos\alpha\end{matrix}\right),$ (3.2)
@@
The stabilizability radius of System 1 is given by
$\tilde{\rho}=|\lambda_{3}|\inf_{l\in\mathbb{N}}\left|\frac{\lambda_{2}}{\lambda_{3}}\frac{\sin(l\alpha-\beta)\pi}{\sin\beta\pi}\right|^{1/(l+1)},$ (3.3)
@@
To reduce the norm of $x$, the general approach is to rotate (apply $J$) as close to $v_{1}$ as possible, then apply $M^{\prime}_{1}$. Consider the switching law in the following general form
$A_{t}=J^{t_{k}}M^{\prime}_{1}J^{t_{k-1}}\cdots M^{\prime}_{1}J^{t_{1}},$ (3.4)
@@
For any vector $x\in\mathbb{R}^{2}$, consider the polar coordinates of $x$, i.e. $x=(r,\;\theta\pi)$, we have
@@
Without losing generality, we assume $v_{1}$ is the $x$ axis, Then we have
@@
Notice that after the first time applying $M^{\prime}_{1}$, the trajectory starts from $v_{2}$, Let $t_{i}=l$ for $i=2\dots k-1$ and $\|M^{\prime}_{1}J^{t_{1}}x\|=C\|x\|$, we have
@@
Since the factor of RHS in (3.7) is independent from $x_{0}$, by taking the infimum over $l$ we obtain the stabilizability radius. ∎
To find the infimum in Lemma 3.1, we need to know $\|l\alpha-\beta\|_{N}$ for $l\in\mathbb{N}$ with $\|\cdot\|_{N}$ from number theory denoting the distance to the closest integer. Now we introduce several lemmas from [1]. The continued fraction of $\alpha$ is given by $\alpha=[a_{0};a_{1},a_{2},\dots]$. Let
@@
and $D_{k}=q_{k}\alpha-p_{k}$, we have the following results.
@@
([1] Lemma (3.1)) For any $\alpha\in\mathbb{R}\setminus\mathbb{Q}$, we have $l\in\mathbb{N}$ satisfy that
@@
where $\delta_{k+1}=c_{k+1}-b_{k+1}$ with $c_{k+1}$ and $b_{k+1}$ from Lemma 3.2 and $3.$ Then we have
@@
Then we have the following theorem.
Theorem 2.
For System 1 with a given $\beta$, we can obtain the stabilizability radius $\tilde{\rho}$ for the following three cases.
@@
    For any $\alpha\in\mathbb{Q}$ and not in case $1$, let $\alpha=p/q$, we have
    $\tilde{\rho}=|\lambda_{3}|\min_{l}\left|\frac{\lambda_{2}}{\lambda_{3}}\frac{\sin(l\alpha-\beta)\pi}{\sin\beta\pi}\right|^{1/(l+1)},\;l\in\{0,\dots,q-1\}$ (3.14)
@@
    $\tilde{\rho}=|\lambda_{3}|\inf_{l_{n}}\left|\frac{\lambda_{2}}{\lambda_{3}}\frac{\sin(l_{n}\alpha-\beta)\pi}{\sin\beta\pi}\right|^{1/(l_{n}+1)}.$ (3.15)
@@
    where $b_{k+1}$ is obtained from Lemma 3.3 by letting $\theta=\beta$ when $\alpha\in(0,1-\beta)$, or $\theta=\beta-1$ when $\alpha\in[1-\beta,1)$.
@@
Case 11 and $2$ are easily obtained from Lemma 3.1. For case 3, choosing
$l_{n}=\sum_{k=0}^{n}b_{k+1}q_{k},$ (3.17)
with $b_{k+1}$ from Lemma 3.3 by letting $\theta=\beta$ when $\alpha\in(0,1-\beta)$, or $\theta=\beta-1$ when $\alpha\in[1-\beta,1)$. Then we have $\delta_{k+1}=0$ for all $k\leq n$. Then $l_{n}\alpha$ is the best approximation of $\beta$ for any $l\leq l_{n}$ by Lemma 3.4, the infimum of $\tilde{\rho}$ reached by one of $l_{n}$ or when $n$ tends to infinity. ∎
@@
The sequence $\{l_{n}\}$ described in Theorem 2 is non-decreasing and unbounded, but not necessarily increasing.
It is worth mentioning that $\tilde{\rho}=\check{\rho}$ in this case, as shown in the proof of Lemma 3.1. Then the lower bound from Theorem 1 gives us a trivial inequality $\tilde{\rho}\geq\tilde{\rho}/2$.
3.2 Zero stability radius
There is one interesting question about System 1: ‘When would $\tilde{\rho}=0$?’ For a given $\beta$, we have the following two cases of $\tilde{\rho}$ for different $\alpha$.
@@ 3.2 Zero stability radius
    The infimum is reached by a finite $l$.
@@
    The infimum is reached when $n$ tends to infinity.
In the first case, $\tilde{\rho}=0$ if and only if $\|l\alpha-\beta\|_{N}=0$ for some integer $l$. Let $L_{\beta}(0)$ denotes the set of $\alpha$ such that the second case holds, it is clear that $L_{\beta}(0)$ only contains irrational $\alpha$, which all belong to the third case in Theorem 2. The Hausdorff dimension of $L_{\beta}(0)$ is given by the following more general lemma.
Lemma 3.5.
For a given $\beta$ and a constant $c\in[0,1)$, let $L_{\beta}(c)$ be the set of $\alpha$ such that $\tilde{\rho}=c$ and the infimum in Lemma 3.1 is reached when $n$ tends to infinity. Then we have that $L_{\beta}(c)$ is a zero Hausdorff dimension set.
@@
with $\{l_{n}\}$ from the third case in Theorem 2. Let $R_{n}$ denote $\|l_{n}\alpha-\beta\|_{N}$, as $\sin x\sim x$ when $x\rightarrow 0$, there exists a infinite subsequence $\{l_{n_{i}}\}$ such that
@@
For any $0<s<1$ and $\epsilon>0$, let $\mathcal{L}(\alpha)$ denotes the smallest integer of each $\alpha$ satisfy that
@@
$\left\|\alpha-\frac{\beta+j}{l_{n_{i}}}\right\|<\frac{c_{1}^{l_{n_{i}}}}{l_{n_{i}}},$ (3.21)
@@
Let $I(k,j)$ be the $j$th interval of $k$ in 3.23, by (d) we have $|I(k,j)|=2c_{1}^{k}/k<\epsilon$. By (c) and (e), we have
@@
Then for any $0<s<1$, the Hausdorff measure of $L_{\beta}(c)$ is zero, $L_{\beta}(c)$ is a zero Hausdorff dimension set. ∎
Theorem 3.
For a given $\beta$, let $S_{\beta}(c)$ be the set of $\alpha$ such that $\tilde{\rho}=c$, then $S_{\beta}(c)$ is a zero Hausdorff dimension set.
@@
We have
$S_{\beta}(c)=\{\alpha|\exists l,\,\|l\alpha-\beta\|_{N}=c\}\cup L_{\beta}(c).$ (3.25)
For any $l\in\mathbb{N}$, there are only finite $\alpha$ satisfies that $\|l\alpha-\beta\|_{N}=0$. Hence $S_{\beta}(c)$ is the union of a countable set and a set with zero Hausdorff dimension. ∎
For rational $\beta$, the $\alpha$ in $L_{\beta}(0)$ are related to a familiar set of numbers.
@@
Let $L$ be the set of all Liouville numbers in $(0,1)$. For rational $\beta$ we have $L_{\beta}(0)\subset L$.

### v3

@@ 3 Rank one singular matrices
In this paper, a matrix set $\mathcal{M}$ is called $\tilde{\rho}-$minimal if for any proper subset $\mathcal{M}_{1}$, i.e. $\mathcal{M}_{1}\subset\mathcal{M}$, we have
@@
Not all linear discrete time switching systems are $\tilde{\rho}-$minimal. However, it is sufficient to only consider the stabilizability radius of $\tilde{\rho}-$minimal matrix sets.
@@
[…]
The proof of Proposition 3.4 trivially comes from Corollary 3.4 in [20]. We have the following proposition of $\mathcal{M}$ with only rank one singular matrices
Proposition 3.5.
For any $n\geq 2$ and $m\in\mathbb{Z}^{+}$, we can find at least one matrix set $\mathcal{M}\subseteq\mathbb{R}^{n\times n}$ with cardinality $m$ such that $\mathcal{M}$ is $\tilde{\rho}-$minimal and contains only rank one singular matrices.
@@
We introduce a $\mathcal{M}$ for $n=2$ and any $m$ in the following. Let $M_{i}$ denote matrices in $\mathcal{M}$, we have
$M_{i}=s_{n}(M_{i})v_{i}w^{T}_{i},\;i\in{1,\dots,m},$ (3.6)
where $s_{n}(M_{i})$ denotes the largest singular value as in section $2$, $v_{i}$ and $w_{i}$ are normalized vectors. Now let $s_{n}(M_{i})=1$ and
$w_{i}=(\cos(\theta_{i}+\epsilon),\sin(\theta_{i}+\epsilon))^{T},\;v_{i}=(\cos(\theta_{i-1}+\pi/2),\sin(\theta_{i-1}+\pi/2))^{T},\;i\in{1,\dots,m},$
where $\theta_{i}\in(0,2\pi)$, for $i\in\{0,\dots,m\}$ and $\epsilon>0$. Then we have
$M_{i}M_{j}=\cos(\theta_{i}-\theta_{j-1}-\pi/2+\epsilon)v_{i}w^{T}_{j},\;i,\;j\in\{1,\dots,m\}.$ (3.7)
Choosing $\theta_{0}=\theta_{m}=3\pi/2$ and all $\theta_{i}$ such that:
$\theta_{i}\neq\theta_{j}+k\pi,\;i\neq j\in\{0,\dots,m-1\},\;k\in\mathbb{Z}.$ (3.8)
Then there exists $\epsilon>0$ such that
$|\sin(\theta_{i}-\theta_{j}+\epsilon)|\geq\sin\epsilon,\;i,j\in\{0,\dots,m-1\}.$ (3.9)
Consider any combination of matrices with length $L$. We have
$A_{\sigma,L}=M_{\sigma(L)}\cdots M_{\sigma(1)}=v_{\sigma(L)}w^{T}_{\sigma(1)}\prod_{i=1}^{L-1}\cos\left(\theta_{\sigma(i)}-\theta_{\sigma(i+1)-1}-\pi/2+\epsilon\right).$ (3.10)
Then the lower spectral radius of $\mathcal{M}$ is given by
$\check{\rho}=\lim_{L\rightarrow\infty}\inf_{\sigma}\|A_{\sigma,L}\|^{1/L}=\lim_{L\rightarrow\infty}\inf_{\sigma}\|v_{\sigma(L)}w^{T}_{\sigma(1)}\|^{1/L}\left|\prod_{i=1}^{L-1}\sin\left(\theta_{\sigma(i)}-\theta_{\sigma(i+1)-1}+\epsilon\right)\right|^{1/L}$ (3.11)
By (3.9) we have
$\check{\rho}\geq\lim_{L\rightarrow\infty}\|v_{\sigma(L)}w^{T}_{\sigma(1)}\|^{1/L}(\sin\epsilon)^{(L-1)/L}\geq\sin\epsilon.$ (3.12)
Choose $\sigma_{k}=m+1-k$. We have
$(M_{1}M_{2}\cdots M_{m})^{k}=(\sin\epsilon)^{mk-1}v_{1}w^{T}_{m},\;k\in\mathbb{Z}^{+},$ (3.13)
is an optimal sequence of $\check{\rho}$. In view of (3.8), equality in (3.9) hold only if $i=j$. Thus the only optimal sequence has $\sigma(i)=\sigma(i+1)-1$ in (3.11), that is, $(M_{1}M_{2}\cdots M_{m})^{k}$. By Proposition 3.1, the optimal sequence of $\tilde{\rho}$ is the same as the optimal sequence of $\check{\rho}$, so this $\mathcal{M}$ is a $\tilde{\rho}-$minimal set. ∎
Let us mention $\mathcal{M}$ with $n=2$ and $m=3$ as an example of the proof of proposition 3.5. Let $v_{1}=(1,0)^{T}$, $\theta_{1}=\pi/6$, $\theta_{2}=5\pi/6$, $w_{3}=(\cos(3\pi/2+\epsilon),\sin(3\pi/2+\epsilon))^{T}$. It is easy to verify that the sequence $M_{i}^{k}$ give us
$\lim_{L\rightarrow\infty}\|A_{\sigma,L}\|^{1/L}=|\cos(\pi/6+\epsilon)|,$
and sequence $(M_{i}M_{j})^{k}$ give us
$\lim_{L\rightarrow\infty}\|A_{\sigma,L}\|^{1/L}=|\sin\epsilon\cos(\pi/6-\epsilon)|^{1/2},$
for $i,j\in\{1,2,3\}$. Finally, the sequence $(M_{1}M_{2}M_{3})^{k}$ give us
$\lim_{L\rightarrow\infty}\|A_{\sigma,L}\|^{1/L}=|\sin\epsilon|.$
Thus for sufficiently small $\epsilon$, $\mathcal{M}$ is a $\tilde{\rho}-$minimal set.
4 Two dimensions: Elliptic case
@@ 4.1 A singular and a rotation matrix
In Corollary 3.2, we mentioned $\tilde{\rho}=\check{\rho}$ for any two switched system with singular matrices. Then to better understand the structure of linear switched systems with singular matrices, we start from two dimensional linear switched systems.
System 4.1.
@@
Let $\lambda_{3}$ and $\overline{\lambda}_{3}$ denote two eigenvalues of $M_{2}$. The matrix $M_{2}$ is similar to its real Jordan form (see for example Theorem 3.4.1.5. in [9]),
$J=|\lambda_{3}|\left(\begin{matrix}\cos\alpha\pi&\sin\alpha\pi\\ -\sin\alpha\pi&\cos\alpha\pi\end{matrix}\right)$
where $\cos{\alpha\pi}={|\lambda_{3}+\overline{\lambda}_{3}|}/2|\lambda_{3}|$. Thus, there exists an invertible matrix $P$ such that $M_{2}=PJP^{-1}$. Let $M^{\prime}_{1}=P^{-1}M_{1}P$, then the system 4.1 is stabilizable if and only if the following switched system is stabilizable.
System 4.2.
@@
$M^{\prime}_{1}=\left(\begin{matrix}a^{\prime}_{11}&a^{\prime}_{12}\\ a^{\prime}_{21}&a^{\prime}_{22}\end{matrix}\right),\quad J=|\lambda_{3}|\left(\begin{matrix}\cos\alpha\pi&\sin\alpha\pi\\ -\sin\alpha\pi&\cos\alpha\pi\end{matrix}\right),$ (4.2)
@@
Let $\lambda_{1}$ and $\lambda_{2}$, $v_{1}$ and $v_{2}$ denote two eigenvalues and normalized eigenvectors of $M^{\prime}_{1}$. If $\lambda_{1}=\lambda_{2}=0$, $M_{1}$ is nilpotent, then System 4.2 becomes trivial. Let $\lambda_{2}\in\mathbb{R}\setminus\{0\}$ and $\lambda_{1}=0$, then we have following results.
@@
The stabilizability radius of System 4.2 is given by
$\tilde{\rho}=|\lambda_{3}|\min\left\{1,\,\inf_{l\in\mathbb{N}}\left|\frac{\lambda_{2}}{\lambda_{3}}\frac{\sin(l\alpha-\beta)\pi}{\sin\beta\pi}\right|^{1/(l+1)}\right\},$ (4.3)
@@
To reduce the norm of $x$, the general approach is to rotate (apply $J$) as close to $v_{1}$ as possible, then apply $M^{\prime}_{1}$. Consider the switching law in following general forms: $A_{\sigma,t}=J^{t}$ or
$A_{\sigma,t}=J^{t_{k}}M^{\prime}_{1}J^{t_{k-1}}\cdots M^{\prime}_{1}J^{t_{1}},$ (4.4)
@@
When $A_{\sigma,t}=J^{t}$, we have $\tilde{\rho}=|\lambda_{3}|$. Now consider the optimal sequence of the form (4.4). For any vector $x\in\mathbb{R}^{2}$, consider the polar coordinates of $x$, i.e. $x=(r,\;\theta\pi)$, we have
@@
Without losing generality, we assume the line in the direction of $v_{1}$ is the $x$ axis, Then we have
@@
Notice that after the first time applying $M^{\prime}_{1}$, the trajectory starts from $v_{2}$. Let $t_{i}=l$ for $i=2\dots k-1$ and $\|M^{\prime}_{1}J^{t_{1}}x\|=C\|x\|$ for some constant $C$ only depending on $t_{1}$ and $x$, and we have
@@
Since $C$ is independent from $k$, from the definition of $\tilde{\rho}$ we have
$\tilde{\rho}=\sup_{x_{0}}\inf_{l}\left|\lambda_{2}|\lambda_{3}|^{l}\frac{\sin(\beta-l\alpha)\pi}{\sin\beta\pi}\right|^{1/(l+1)}=\inf_{l}\left|\lambda_{2}|\lambda_{3}|^{l}\frac{\sin(\beta-l\alpha)\pi}{\sin\beta\pi}\right|^{1/(l+1)}.$ (4.8)
To find the infimum in Lemma 4.3, we need to know $\|l\alpha-\beta\|_{N}$ for $l\in\mathbb{N}$ with $\|\cdot\|_{N}$ from number theory denoting the distance to the closest integer. This is a Diophantine approximation problem and we need the continued fraction of $\alpha$ to solve it(see for example [18]). Let $a_{0};a_{1},a_{2},\dots]$ be the continued fraction of $\alpha$,
@@
and $D_{k}=q_{k}\alpha-p_{k}$. Knowing these, let us introduce several lemmas from [1].
@@
([1] Lemma (3.1)) For any $\alpha\in\mathbb{R}\setminus\mathbb{Q}$, we have $l\in\mathbb{N}$ satisfies
@@
where $\delta_{k+1}=c_{k+1}-b_{k+1}$ with $c_{k+1}$ and $b_{k+1}$ from Lemma 4.4 and 4.5. Then we have
@@
Then we have more information about the infimum in Lemma 4.3.
Theorem 4.7.
For System 4.2 with a given $\beta$, we can obtain the stabilizability radius $\tilde{\rho}$ for the following three cases.
@@
    For any $\alpha\in(0,1)\cap\mathbb{Q}$ and not in case $1$, let $\alpha=p/q$, then we have
    $\tilde{\rho}=|\lambda_{3}|\min\left\{1,\,\min_{l}\left|\frac{\lambda_{2}}{\lambda_{3}}\frac{\sin(l\alpha-\beta)\pi}{\sin\beta\pi}\right|^{1/(l+1)}\right\},\;l\in\{0,\dots,q-1\}$ (4.15)
@@
    $\tilde{\rho}=|\lambda_{3}|\inf_{n}\left|\frac{\lambda_{2}}{\lambda_{3}}\frac{\sin(l_{n}\alpha-\beta)\pi}{\sin\beta\pi}\right|^{1/(l_{n}+1)}.$ (4.16)
@@
    where $b_{k+1}$ is obtained from Lemma 4.5 by letting $\theta=\beta$ when $\alpha\in(0,1-\beta)$, or $\theta=\beta-1$ when $\alpha\in[1-\beta,1)$. In this case, System 4.2 is stabilizable if $|\lambda_{3}|\leq 1$.
@@
For any $\alpha$ such that $\|l\alpha-\beta\|_{N}=0$, we have $\sin\|l\alpha-\beta\|_{N}\pi=0$, then by lemma 4.3 we have $\tilde{\rho}=0$. For any rational $\alpha$ and not in case $1$, the set $\{\|l\alpha\|_{N}|l\in\mathbb{N}\}$ is a finite set, and if we let $\alpha=p/q$, then
$\|(l+q)\alpha\|_{N}=\|l\alpha\|_{N}.$ (4.18)
then by Lemma 4.3, the infimum of $\tilde{\rho}$ is $|\lambda_{3}|$ or a minimum and reached by one of $l\in\{0,\dots,q-1\}$.
Now let us prove the case 3. Let $\theta=\beta$ when $\alpha\in(0,1-\beta)$, and $\theta=\beta-1$ when $\alpha\in[1-\beta,1)$. From Lemma 4.5, we know that $\theta=\sum_{k=0}^{\infty}b_{k+1}D_{k}$. Consider minimizing $\|l\alpha-\theta\|$ by a bounded $l$ first, then we only need $k_{0}$ in Lemma 4.6 as large as possible. By choosing
$l_{n}=\sum_{k=0}^{n}b_{k+1}q_{k},\;n\in\mathbb{N},$ (4.19)
we can guarantee that $k_{0}=n+1$ and $\delta_{k+1}=-b_{k+1}$ for $k\geq k_{0}$. This means that $l_{n}\alpha$ is the best approximation of $\beta$ for any $l<l_{n+1}$, i.e. $\|l_{n}\alpha-\beta\|\leq\|l\alpha-\beta\|$ for $l<l_{n+1}$. Now we consider the effect of power term $1/(l_{n}+1)$. As the set $\{\|l\alpha\|_{N}\}$ is dense in $(0,1)$, then there exist an $l_{0}$ such that
$\left|\frac{\lambda_{2}}{\lambda_{3}}\frac{\sin(l_{0}\alpha-\beta)\pi}{\sin\beta\pi}\right|<1.$ (4.20)
Moreover, for $l_{n}\geq l_{0}$ defined above, we always have
$\left|\frac{\lambda_{2}}{\lambda_{3}}\frac{\sin(l_{n}\alpha-\beta)\pi}{\sin\beta\pi}\right|^{1/(l_{n}+1)}<\left|\frac{\lambda_{2}}{\lambda_{3}}\frac{\sin(l\alpha-\beta)\pi}{\sin\beta\pi}\right|^{1/(l+1)},$ (4.21)
for $l_{n}<l<l_{n+1}$. The infimum is reached by one of the $l_{n}$ or when $n\rightarrow\infty$. We also know $\tilde{\rho}<1$ if $\lambda_{3}\leq 1$ from Lemma 4.3 and (4.20) ∎
@@
The sequence $\{l_{n}\}$ described in Theorem 4.7 is non-decreasing and unbounded, but not necessarily increasing.
Remark.
The infimum is often reached by a finite $l_{n}$ in case 3. We will discuss this more in the next subsection.
4.2 Finiteness property
For a given $\beta$, we have the following two cases of $\tilde{\rho}$ for different $\alpha$.
@@ 4.2 Finiteness property
    The infimum in Lemma 4.3 is reached by a finite $l$.
@@
    The infimum in Lemma 4.3 is reached when $n$ tends to infinity.
In the literature, we have a name for case $1$:
Definition 4.8.
A system has finiteness property w.r.t. $\tilde{\rho}$ if there exists a finite matrix sequence $A_{\sigma,t}=M_{\sigma(t)}\cdots M_{\sigma(1)}$ such that $\tilde{\rho}^{t}$ is the spectral radius of $A_{\sigma,t}$.
This finiteness property w.r.t. $\hat{\rho}$ has many known results ([10], for example). We are interested in the finiteness property w.r.t. $\tilde{\rho}$ for system 4.1.
Let $L_{\beta}(c)$ denote the set of $\alpha$ such that the second case holds and $\tilde{\rho}=c$ for some constant $c$. It is clear that $L_{\beta}(c)$ only contains irrational $\alpha$, which all belong to the third case in Theorem 4.7. The dimension of some $L_{\beta}(c)$ are given by the following Theorem.
Theorem 4.9.
The parameter set $L_{\beta}(c)$ for $0\leq c<1$ has zero Hausdorff dimension.
@@
with $\{l_{n}\}$ from the third case in Theorem 4.7. Let $R_{n}$ denote $\|l_{n}\alpha-\beta\|_{N}$, as $\sin x\sim x$ when $x\rightarrow 0$, there exists a infinite subsequence $\{l_{n_{i}}\}$ such that
@@
For any $0<s<1$ and $\epsilon>0$, let $\mathcal{L}(\alpha)$ denote the smallest integer of each $\alpha$ that satisfies
@@
$\left\|\alpha-\frac{\beta+j}{l_{n_{i}}}\right\|_{N}<\frac{c_{1}^{l_{n_{i}}}}{l_{n_{i}}},$ (4.25)
@@
Let $I(k,j)$ be the $j$th interval of $k$ in 4.27, by (d) we have $|I(k,j)|=2c_{1}^{k}/k<\epsilon$. By (c) and (e), we have
@@
Then for any $0<s<1$, the Hausdorff measure of $L_{\beta}(c)$ is zero, so $L_{\beta}(c)$ is a zero Hausdorff dimension set. ∎
We have a immediate corollary from Theorem 4.9.
Corollary 4.10.
For a given $\beta$, $c\in[0,1)$, let $S_{\beta}(c)$ be the set of $\alpha$ such that $\tilde{\rho}=c$, then $S_{\beta}(c)$ is a zero Hausdorff dimension set.
@@
Since $0\leq c<1$, from Lemma 4.3, we have
$S_{\beta}(c)=L_{\beta}(c)\bigcup\left\{\alpha\,|\,\exists l,\,\|l\alpha-\beta\|_{N}\pi=\arcsin\left|\frac{c^{l+1}\sin\beta\pi}{\lambda_{2}\lambda^{l}_{3}}\right|\right\}.$ (4.29)
For any $l\in\mathbb{N}$, there are only finite $\alpha$ satisfies that $\|l\alpha-\beta\|_{N}$ equals to a known constant. Then $S_{\beta}(c)$ is the union of a countable set and a set with zero Hausdorff dimension, hence a zero Hausdorff dimensional set. ∎
We know that System 4.2 has finiteness property w.r.t. $\tilde{\rho}$ for $\alpha\in(0,1)\setminus\cup_{c}L_{\beta}(c)$. However, This union of $L_{\beta}(c)$ over $c$ is not necessarily a zero Hausdorff dimensional set, so we need more restriction on System 4.2 to ensure it has zero Hausdorff dimension.
Corollary 4.11.
For System 4.1 with fixed $\beta$, $\cup_{c}L_{\beta}(c)\cap S$ has zero Hausdorff dimension if $|\lambda_{3}|\leq 1$, where $S$ is denoted by
$S=\left\{\alpha\in(0,1)\;:\;\inf_{l}\left|\frac{\lambda_{2}}{\lambda_{3}}\frac{\sin(l\alpha-\beta)\pi}{\sin\beta\pi}\right|^{1/(l+1)}<c_{0}<1\right\},$ (4.30)
for an arbitrary constant $c_{0}\in(0,1)$.
Proof.
As mentioned in the proof of Theorem 4.7, $\{l\alpha\}$ is dense in $(0,1)$ for irrational $\alpha$, so $S$ is not empty. Consider System $1$ with $|\lambda_{3}|\leq 1$. By lemma 4.3, we know $S$ is equivalent to the set
$S=\{\alpha\in(0,1)|\tilde{\rho}<c_{0}<1\}.$ (4.31)
Then for any $\alpha\in\cup_{c}L_{\beta}(c)\cap S$, we have $R_{n_{i}}=o(c_{0}^{l_{n_{i}}})$, where $R_{n_{i}}$ is same as in the proof of Theorem 4.9. Similar to the proof of Theorem 4.3, for any $0<s<1$ and $\epsilon>0$, let $\mathcal{L}(\alpha)$ denotes the smallest integer of each $\alpha$ satisfy that
  • $\mathcal{L}(\alpha)\in\{l_{n_{i}}\}$,
  • $\|\mathcal{L}(\alpha)\alpha-\beta\|_{N}<c_{0}^{\mathcal{L}(\alpha)}$,
  • $\left(t^{1-s}c_{0}^{st}\right)^{\prime}|_{t=(\mathcal{L}(\alpha)-1)}<0$
  • $2c_{0}^{\mathcal{L}(\alpha)}/\mathcal{L}(\alpha)<\epsilon$,
  • $2^{s}(-s\ln c_{0})^{s-2}\Gamma(2-s,-s\ln c_{0}(\mathcal{L}(\alpha)-1))<\epsilon$,
and $\mathcal{L}=\inf\{\mathcal{L}(\alpha)|\alpha\in\cup_{c}L_{\beta}(c)\cap S\}$. Then there exists an $i$ such that $l_{n_{i}}>\mathcal{L}$ and
$\alpha\in\left(\frac{\beta+j}{l_{n_{i}}}-\frac{c_{0}^{l_{n_{i}}}}{l_{n_{i}}},\frac{\beta+j}{l_{n_{i}}}+\frac{c_{0}^{l_{n_{i}}}}{l_{n_{i}}}\right).$ (4.32)
We have then constructed a cover of $\cup_{c}L_{\beta}(c)\cap S$ given by
$\bigcup_{c}L_{\beta}(c)\cap S\subseteq\bigcup_{k=\mathcal{L}}^{\infty}\bigcup^{k-1}_{j=0}\left(\frac{\beta+j}{l_{n_{i}}}-\frac{c_{0}^{l_{n_{i}}}}{l_{n_{i}}},\frac{\beta+j}{l_{n_{i}}}+\frac{c_{0}^{l_{n_{i}}}}{l_{n_{i}}}\right),$ (4.33)
and same as the proof of Theorem 4.9, each interval and the union give zero Hausdorff measure for $\cup_{c}L_{\beta}(c)\cap S$. ∎
From 4.11, we know that if we only consider System 4.2 with a given $\beta$, $\alpha\in S$ and $|\lambda_{3}|\leq 1$. Then the subset of $S$, such that System 4.2 does not have finiteness property w.r.t. $\tilde{\rho}$, is a zero Hausdorff dimensional set.
For rational $\beta$, we do not need to consider $S$.
@@
For rational $\beta$ we have $\cup_{c\in[0,1)}L_{\beta}(c)\subset L$, where $L$ denotes the set of all Liouville numbers in $(0,1)$.